OpenStudy (anonymous):

A gas at 407 K occupies 6.65 L at a pressure of 1.98 atm. What will its pressure (in atm) be at 249 K if its volume is reduced to 4.30 L? Report your answer to 3 significant figures

OpenStudy (anonymous):

ok i keep repeating this but ill keep repeating it... so we have here a closed system and the amount of gas is same and only other conditions are changed!!! so we start with basic equation: pV=nRT where now on 1 is index of conditions at the begining and 2 is index of conditions at the end (after change)! if amount isnt changed we can now write: n1=n2 n1=p1V1/RT1 n2=p2V2/RT2 so we have: p1V1/RT1= p2V2/RT2 cause we have R at both sides and at same positions we can cross them out and now we have: p1V1/T1 = p2V2/T2 now we need to get p2 so we do some math magic and we get: p2= p1V1T2/V2T1